hali went shopping and spent half of her money on an ipod. THEN SHE SPENT A QUARTER OF WHAT SHE HAD LEFT ON A GRILL SET.tHEN SHE SPENT A FIFTH OF HER MONEY REMAINING ON A SWEATER AND THEN SHE BOUGHT TWO BOARD GAMES WITH ATHIRD OF WHAT SHE HAD REMAINING AFTER THAT. THEN SHE SPENT ANOTHER $50 ON LOTION. aFTER THAT SHE SPENT $18.50 ON LUNCH. SHE LAFT THE MALL WITH $41.50.
I added 41.50 and 18.50 and got 60 then i added 50 to that and got 110, now I don't know what to do next.

Answer by stanbon!(97)

Let's keep track of how much is left after each transaction,
rather than how much was spent.
Let's say Hali started with amount x of money.
Ipod took 1/2 of it leaving her with (1/2)x
Grill cost her 1/4 of that leaving her with (3/4)(1/2)x = (3/8)x
Sweater cost 1/5 of that leaving her (4/5)(3/8)x= (3/10)x
Games cost 1/3 of that leaving her (2/3)(3/10)x = (1/5)x
Lotion and Lunch cost her 68.50 leaving her (1/5)x-68.50
EQUATION:
(1/5)x-68.50 = 41.50
(1/5)x = 110
x = $550.00 (the amount she started with)
Cheers,
Stan H.